OpenStudy (anonymous):

Find the summation of the finite geometric series below. (It should equal 3.75)

OpenStudy (anonymous):

\[\sum_{i = 1}^{10}5(\frac{ -1 }{ 3 })^{i-1}\]

OpenStudy (anonymous):

Through the equation, I keep getting 1.5

jimthompson5910 (jim_thompson5910):

In this case, the first term is a = 5 and the common ratio is r = -1/3

jimthompson5910 (jim_thompson5910):

you would then plug these values into the nth partial sum formula \[\Large S_n = a\left(\frac{1-r^n}{1-r}\right)\]

jimthompson5910 (jim_thompson5910):

oh and in this case, n = 10 since you want the 10th partial sum

OpenStudy (anonymous):

I got it now, thanks.

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

I think I was multiplying it by 2 for some reason